OpenStudy (ttp):

Use the formula A=lw Find A when I=8 And W =12.1

OpenStudy (ttp):

Show your work as you solve the equation

OpenStudy (ttp):

@e.cociuba

OpenStudy (e.cociuba):

Use the formula A=lw Find A when I=8 And W =12.1 Plug in your numbers into the formula and solve.

OpenStudy (e.cociuba):

Multiply "I" by "w".

OpenStudy (ttp):

after i multiply i by w what do i do?

OpenStudy (anonymous):

Thats it. You have the answer

OpenStudy (e.cociuba):

After you multiply, you get your answer. A=?

OpenStudy (e.cociuba):

You're trying to find out what "A" equals.

OpenStudy (ttp):

So 12.1 * 8 is my equation?

OpenStudy (anonymous):

A=LW A = 8 * 12.1 A = 96.8

OpenStudy (anonymous):

No, 8 * 12.1 is. :)

OpenStudy (e.cociuba):

@Niall did it correct as shown above.

OpenStudy (ttp):

Question where did you get tte point 8 from? :) I know how you got 96

OpenStudy (anonymous):

Because your multiplying by 12.1, not 12. So 8 * 12.1 = 96.8

OpenStudy (ttp):

Thank y'all so muchh!!

OpenStudy (anonymous):

Your Welcome! :)

OpenStudy (e.cociuba):

Use the formula A=lw Find A when I=8 And W =12.1 Plug in your numbers that they list above in the formula. So, A=Iw I=8 w=12.1 Multiply to get A.